OpenStudy (anonymous):

If x^2 = 20, what is the value of x? ±square root of 10 ±square root of 20 ±10 ±40

OpenStudy (mertsj):

\[x^2=20\] \[\sqrt{x^2}=\pm \sqrt{20}\] \[x=\pm \sqrt{20}\]

OpenStudy (anonymous):

I have no clue what the answer is but this website will help you out with it and most likly give you the answer to it. The only reason why i am doing this is cause you do need to learn it too. http://www.wou.edu/~brownk/Math095/Math095.09.4.SquareRootPropertyEquations.f.pdf